OpenStudy (anonymous):

So lost.. NEED help!!! Part 1: Provide the solution to the equation 25 = √5 ^ (2x + 6). Part 2: Using complete sentences, explain the strategy used in solving this equation the way you did. I just need help with part 2 now

OpenStudy (anonymous):

do you got a map?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

I just need help with part 2 now

OpenStudy (phi):

If you got the answer, just write down how you got it. That is the answer to part 2.

OpenStudy (anonymous):

I don't know how to get it. The other guy gave it to me @phi

OpenStudy (phi):

do you know \( \sqrt{5}= 5^{\frac{1}{2}} \). They mean the same thing. do you know 25= 5*5 or 5^2 The first thing you should do is rewrite the equation using 5^power instead of 25 and sqrt(5)